OpenStudy (anonymous):

Simplify. Express the quotient as a radical expression. 36^ 1/2

OpenStudy (anonymous):

=6

OpenStudy (anonymous):

would that be -6

OpenStudy (anonymous):

=+/- 6

OpenStudy (anonymous):

no, it's just sqrt(36)=6

OpenStudy (anonymous):

no +, - , that would be onle when we have unknowns

OpenStudy (anonymous):

yeah but they mutiple choice answers only gives -6

OpenStudy (anonymous):

like x^2 = 36^(1/2) ===> x= + - 6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

awesome thanks

OpenStudy (anonymous):

you are welcome.